OpenStudy (anonymous):

Number of points if discontinuity of [6sinx] x belongs to [o,pi] and [.] denotes greatest integer function

OpenStudy (anonymous):

On \(0\le x\le\pi\) we find \(0\le\sin x\le 1\) and thus \(0\le6\sin x\le6\)

OpenStudy (anonymous):

Cant understand

OpenStudy (anonymous):

Ok clr

OpenStudy (anonymous):

Nxt¿

OpenStudy (anonymous):

Note that our function is essentially identical apart from reversal on \([0,\pi/2],[\pi/2,\pi]\). On each we see a monotone progression from \(0\) to \(6\) (and vice versa) and thus \(5+1+5=11\) total discontinuities

OpenStudy (anonymous):